OpenStudy (anonymous):

Evaluate x*2-(y-2) squared when y=3 and x=3? I got seven, but I don't believe that is right.

OpenStudy (anonymous):

\[x^2-(y-2)\] replace and get \[3^2-(3-2)=9-1=8\]

OpenStudy (anonymous):

do math in this order: brackets > multiply and divide > add and minus

OpenStudy (anonymous):

oh wait, is it \[x^2-(y-2)^2\]?

OpenStudy (anonymous):

no matter it is 8 in that case too

OpenStudy (anonymous):

it's\[2x - (y-2)\]